# Novel Dynamical Systems with Finite-Time and Fixed-Time Stability for Generalized Inverse Mixed Variational Inequality Problems

### v1

@@ 6 Numerical illustration
We run Algorithm (17) with parameters $\gamma=1$, $a_{1}=0.9$, $a_{2}=0.5$, $a_{3}=0.0001$, $k_{1}=0.4$, $k_{2}=1.5$, and $\theta=0.0001+\frac{1}{k}$, where $k$ denotes the number of iteration. Algorithm (14) is implemented with parameters $k=3$, $\tau=1$, and $\theta=0.2$. Starting from the initial value $w_{0}=50.0$, after $200$ iterations the sequence generated by (17) converges to the approximate solution $3.3339186877\times 10^{-5}$, whereas the sequence generated by Algorithm (14) converges to the approximate solution $1.8750000000\times 10^{-3}$.
@@
By Remark 4, the dynamical system in [16] is a special case of (8)–(9) when $F$ is the identity mapping, and its numerical examples therefore illustrate the effectiveness of our proposed algorithm in this case. Moreover, when $a_{3}=0$, and $F$ is the identity mapping, (8)–(9) reduces to the system studied in [10], so the numerical examples therein can likewise be used to demonstrate the performance of our method.
@@ 7 Conclusion
This work introduces two novel dynamical-system-based approaches for solving generalized inverse mixed variational inequality problems. The first system exhibits finite-time stability, while the second achieves fixed-time stability. In particular, the fixed-time stable model guarantees the existence and uniqueness of the solution, as well as convergence of the associated trajectories to the unique solution of the corresponding the GIMVIP within a uniform time bound. It is further demonstrated that this fixed-time convergence property is retained when the continuous-time dynamics are discretized via a forward Euler scheme. Numerical simulations are conducted to confirm the effectiveness of the proposed algorithm.
Possible directions for future research include extending the finite-time and fixed-time stability analysis of the proposed projection-type dynamical systems to more general frameworks, such as infinite-dimensional Hilbert or Banach spaces, and relaxing the imposed assumptions on the underlying operators and model parameters.

### v2

@@ 6 Numerical illustration
Algorithm (23) is implemented with the parameters $\gamma=1$, $a_{1}=0.9$, $a_{2}=0.5$, $a_{3}=10^{-4},G_{d}=T_{d}=1$, $k_{1}=0.4$, $k_{2}=1.5$, and $\theta=10^{-4}+\frac{1}{n}$, where $n$ denotes the iteration index. Algorithm (20) is implemented with the parameters $\tau=1$ and $\theta=0.2$.
Starting from the initial value $w_{0}=50.0$, after 150150 iterations the sequence generated by (23) converges to the approximate solution $5.33\times 10^{-5}$ when $k_{3}=1$, and to $1.13\times 10^{-4}$ when $k_{3}=0$. In contrast, the sequence generated by Algorithm (20) converges to the approximate value $2.08$ for $k=3$, and to $1.39\times 10^{-3}$ for $k=2$.
@@
By Remark 5, the dynamical system proposed in [17] is a special case of (8) when $F$ is the identity mapping and $k_{3}=1$. Consequently, the numerical examples presented in [17] illustrate the effectiveness of our proposed algorithm in this setting. Likewise, the differential equation system studied in [31] coincides with (8) when $F$ is the identity mapping and $k_{3}=0$; hence, the numerical examples therein can also be used to demonstrate the performance of our method. Moreover, when $a_{3}=0$ and $F$ is the identity mapping, system (8) reduces to the one investigated in [11], and the corresponding numerical examples further validate the effectiveness of our approach.
@@ 7 Conclusion
This work has developed three dynamical-system-based approaches for solving generalized inverse mixed variational inequality problems. Specifically, we proposed continuous-time dynamical systems that achieve finite-time, fixed-time, and predefined-time stability, respectively. The finite-time stable system ensures convergence within a finite settling time that depends on the initial condition, whereas the fixed-time stable system guarantees convergence within a uniform time bound independent of the initial state. More importantly, the predefined-time stable system allows the convergence time to be explicitly prescribed in advance through appropriate parameter selection, providing greater flexibility and user autonomy in time-critical applications.
For the fixed-time and predefined-time stable models, we established the existence and uniqueness of solutions to the corresponding GIMVIPs and proved that the generated trajectories converge to the unique solution within the specified time bounds. Furthermore, it was shown that the convergence properties of the continuous-time dynamics are preserved under an explicit forward Euler discretization, leading to a proximal point-type iterative algorithm with guaranteed fixed-time and predefined-time convergence. Numerical experiments were presented to validate the theoretical findings and to demonstrate the effectiveness and accelerated convergence behavior of the proposed methods.
Several directions for future research merit further investigation. These include extending the proposed finite-time, fixed-time, and predefined-time stability analyses to more general settings, such as infinite-dimensional Hilbert or Banach spaces, as well as relaxing the assumptions imposed on the underlying operators and model parameters. Another promising avenue is the development of adaptive or data-driven parameter selection strategies for predefined-time stable algorithms, which could further enhance their applicability in large-scale optimization and inverse equilibrium problems.